OpenStudy (anonymous):

find the fixed point of the function f(x)=x^2-4x+6

OpenStudy (amistre64):

??? you mean the vertex?

OpenStudy (turingtest):

no it is when f(c)=c

OpenStudy (anonymous):

It just says the fixed point or points oif the function. I think that is why I am confused

OpenStudy (amistre64):

subtract an x and quad formula then if thats the case ;)

OpenStudy (turingtest):

This can be handled easily with the quadratic equation it would seem.

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

I.e., as TT points out, what are the numbers c, if any such that f(c) = c Solve that equation.

OpenStudy (anonymous):

where are you getting the f(c)=c

OpenStudy (jamesj):

This is the definition of a fixed point.

OpenStudy (anonymous):

oh they didn't tell me that. That would have been nice for my instructor to tell me that. Just starting to learn this concept.

OpenStudy (jamesj):

Call it c, p, x, whatever you like. A number c in the domain of a function f is called a FIXED POINT of f if it has the property that f(c) = c.

OpenStudy (jamesj):

The intuition for the name is clear: the function f leaves fixed points where they are; they are not transformed to other values.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you walk me through this

OpenStudy (anonymous):

Does fixed point mean: x^2-4x+6 = x

OpenStudy (anonymous):

That is what I was going to ask

OpenStudy (jamesj):

Geometrically, this is where the graph of f intersects the straight line y = x. So by definition, a fixed point satisfies the equation f(c) = c. That is c^2-4c+6 = c Now determine what values, if any, of c satisfy that equation.

OpenStudy (anonymous):

x^2-4x+6 = x x^2 - 5x + 6 = 0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Then solve for x

OpenStudy (anonymous):

(x-2)(x-3) = x^2 - 5x + 6 x = 2 x = 3 You get points (2,2) (3,3)

OpenStudy (jamesj):

here it is geometrically,

OpenStudy (anonymous):

thanks once I seen it draw out that seems to help.